Wednesday, December 07, 2005

Optical Path Alignment

I placed order of the 10k Ohm 225W resistor this morning, it should be received tomorrow morning. So I cannot run the Prometheus today. I checked the other part of the charging resistors, it looks good except some connectors. I fixed the connectors later. Mike helped me to find some resistors, unfortunately they were 100k Ohms.

In the afternoon, I aligned the optical path of the front end. The main point is the telescope alignment, it took me several hours to send the laser beam straightly into the Prometheus machine. I also found the first 2 mirrors not damaged, but the last one or two mirrors not good. I could see the shadows from them, it means the structure was broken inside.

No comments: